A 100 kg block is pulled at a constant speed of 5 m/s across a horizontal force by an applied force of 122 N directed 37 o above the horizontal. The rate at which the force does work on the block is.


1
Expert's answer
2022-02-21T12:05:54-0500

Given:

"m=100\:\rm kg"

"v=5\:\rm m/s"

"\theta=37^\circ"

"F=122\:\rm N"


The rate of work done (power) is given by

"P=\frac{W}{t}=\frac{Fd\cos\theta}{t}=Fv\cos\theta"

"P=122*5*\cos37^\circ=487\:\rm W"